OpenStudy (peterg2001):

help will medal Alice solved the following equation: -18 - 3x = 12 -3x = 6 x =-2 Which of the following statements is true? Alice is correct. The 6 should be 30 in step two. The 6 should be -6 in step two. The 6 should be -4 in step two.

OpenStudy (peterg2001):

@shinebrightlikeadimon

OpenStudy (anonymous):

B

OpenStudy (anonymous):

-18 - 3x = 12 +18 +18 -3 x = 30 /-3 /-3 x = -10

OpenStudy (anonymous):